OpenStudy (anonymous):

In a quadratic equation, how do you know when to change the outside number to a positive?

OpenStudy (anonymous):

give an example

OpenStudy (anonymous):

Like when you plug b into the formula outside of the square root sign, you're supposed to change it to a negative, but sometimes the answers are positive? Hopefully that makes sense :P

OpenStudy (amistre64):

include the operator

OpenStudy (amistre64):

in general: ax^2 + bx + c is a quadratic equation, but b can be negative ax^2 + (-b)x + c = ax^2 -bx + c

OpenStudy (amistre64):

x^2 - 4x + 7 b = -4 --------------- 2x^2 +2x - 1 b = +2 ---------------- 1/3 x^2 - 4/5 x + 3 b = -4/5 etc ....